# MambaTAD: When State-Space Models Meet Long-Range Temporal Action Detection

### v1

@@ III-D Building Projection Layers
The blocks designed in previous TAD networks [zhang2022actionformer, shi2023tridet, yang2024dyfadet] mainly follow the $Norm\rightarrow Attention/CNN\rightarrow Norm\rightarrow MLP$ flow in each pyramid layer. To address challenges related to computational complexity [DBLP:conf/icml/ZhuL0W0W24], feature discrimination [shi2023tridet, yang2024dyfadet], and receptive field limitations [luo2016understanding, guo2024mambair], we employ Mamba [gu2023mamba] to model global dependencies. However, in Section III-C, we analyze that the vanilla Mamba’s causality mechanism leads to a loss of temporal information, making it unsuitable for the TAD task. This underscores the potential of the proposed DMBSS block in effectively capturing action features for improved localization. The experimental results in Table VII in Section IV-F further validate our analysis.
The DMBSS projection layer. In this section, we construct projection layers utilizing the DMBSS discussed previously. As shown in Fig. 2, each projection layer consists of a DMBSS and a Max pooling layer [tang2023temporalmaxer], because we aim to preserve the benefits of long sequence modeling while enhancing the discriminability of temporal features, all within the linear computational complexity. The two blocks are given by:
@@ III-E Global Feature Fusion Head
The detection head plays a pivotal role in TAD, as it is directly linked to the final prediction outcome. Inspired by [yu2024mambaout, zhu2024temporally], it is evident that the visual state-space model excels in tasks characterized by long sequences. Consequently, we innovatively concatenate features from different levels of the pyramid sequentially to create an extended sequence:
@@
where $\mathbf{F}\in\mathbb{R}^{b\times c\times(s+{s\over 2}+\ldots+{s\over 2^{l-1}})}$, and $l$ is the layer of pyramid. To address the lack of global awareness caused by hierarchical feature structure [lin2017feature, liu2018path], a global feature fusion mechanism is designed to ensure the full use of long-range information across different layers and improve regression and classification accuracy. To this end, we propose the residual global level detection head to adapt DMBSS for action detection. As shown in Fig. 3, we first use the LayerNorm (LN) followed by the DMBSS to capture the temporal long-range dependencies. Then, a residual connection is utilized for easier optimization. Lastly, we use a successive LN to ensure stable feature distribution. The final output can be mathematically represented as follows:
@@
By concatenating features from different pyramid levels into an extended sequence, our design enables the detection head to access temporal information across multiple granularities simultaneously. Traditional hierarchical feature pyramid approaches [lin2017feature, zhang2022actionformer] process each level independently, which limits the receptive field and reduces global context. In contrast, the extended sequence allows DMBSS to operate over a longer, contiguous feature stream, capturing temporal long-range dependencies across all pyramid layers. The residual global feature fusion further stabilizes the feature distribution and facilitates optimization. Together, this design enhances global awareness and improves both classification and boundary regression accuracy in temporal action detection.
@@ III-F State-Space Temporal Adapter
In Sections III-D and III-E, we construct the action detector with projection layers and a global feature fusion head. End-to-end TAD demands strong temporal modeling to bridge low-level encoding and high-level action detection. However, fine-tuning the entire backbone is costly [liu2024end] and may hinder generalization. On the other hand, it remains unexplored whether Mamba can serve as an adapter to bridge large pre-trained foundational models[tong2022videomae, wang2023videomae, wang2024internvideo2] and downstream video understanding tasks. To address these, we propose SSTA, a lightweight yet expressive mechanism for refining temporal representations efficiently.
The standard adapter [houlsby2019parameter] is widely used in PEFT but focuses on the channel dimension, limiting its ability to capture temporal dependencies. AdaTAD [liu2024end] addresses this with the Temporal-Informative Adapter (TIA), combining temporal depth-wise convolution and a fully connected layer. However, depth-wise convolution constrains the receptive field [luo2016understanding, ding2022scaling], while the fully connected layer increases trainable parameters. To overcome these challenges, we propose the State-Space Temporal Adapter (SSTA), as illustrated in Fig. 2. SSTA integrates the DMBSS to enhance temporal modeling. This design enables bidirectional feature interaction while maintaining controlled diagonal masking, effectively capturing both long-range dependencies and local discriminative patterns. The formulation of SSTA is given by:
@@
where $\mathbf{x},\mathbf{x}^{\prime\prime}\in\mathbb{R}^{t\times h\times w\times d}$ denote input and output features, $t$ represents the number of frames, $h$ and $w$ correspond to the spatial dimensions of the video, and $d$ is the hidden feature dimensionality. The function $DMBSS(\cdot)$ represents our proposed block. Following AdaTAD[liu2024end], we employ GELU [hendrycks2016gaussian] as the non-linear activation function, denoted as $\theta(\cdot)$. The transformation matrices $\mathbf{W}^{T}_{down}\in\mathbb{R}^{d\times\frac{d}{\lambda}}$ and $\mathbf{W}^{T}_{up}\in\mathbb{R}^{\frac{d}{\lambda}\times d}$ are trainable projection weights, where $\lambda$ is a scalable factor which controls intermediate dimensions. $\lambda>1$ is enforced to ensure parameter efficiency and reduction. One residual connection is used for better convergence.
@@ IV-A Experimental Setup
Dataset. We experiment on five challenging datasets: THUMOS14 [thumos14], ActivityNet-1.3 [caba2015activitynet], MultiThumos[yeung2018every], HACS-Segment [zhao2019hacs], and FineAction [liu2022fineaction]. The THUMOS14 dataset includes 20 sports action categories, with 200 training and 213 testing untrimmed videos, covering 3,007 and 3,358 action instances. ActivityNet-1.3 features 200 categories across approximately 20K videos and 600 hours, divided into 10K training, 4,926 validation, and 5,044 test videos. MultiThumos introduces more complex scenarios, presenting a greater challenge for TAD. It extends the THUMOS14 dataset with dense, multilabel, frame-level action annotations spanning 30 hours across 400 videos. The dataset comprises 38,690 annotations covering 65 action classes, with an average of 1.5 labels per frame and 10.5 action classes per video, significantly increasing the task’s difficulty. HACS-Segment has 200 categories, with 37.6K training and 5,981 test videos. FineAction offers 103K temporal instances across 106 fine-grained actions in 17K untrimmed videos, with 8,440 training, 4,174 validation, and 4,118 testing. Only training and validation annotations are available. We adhere to standard protocols for training and evaluation.
@@ IV-B Comparison With the State of the Art Methods
THUMOS14 & ActivityNet-1.3 with off-the-shelf features. Firstly, we compare MambaTAD with leading TAD methods on THUMOS14 and ActivityNet-1.3, as shown in Table I. Typically, TAD models achieve better performance with advanced backbones and optical flow, as evidenced by the second part of the table. Our evaluation encompasses a wide range of models, including CNNs, GCNs, Transformer-based, Diffusion-based, and Mamba-based architectures. For a fair comparison, we provide the results of MambaTAD using traditional features in the first part, highlighting its strong adaptability and compatibility. Specifically, MambaTAD surpasses the previous SOTA by 0.5% on THUMOS14 with I3D [carreira2017quo] features and 1.2% on ActivityNet-1.3 using R(2+1)D [alwassel2021tsp] features. In the second part, MambaTAD achieves an average mAP of 73.9% on THUMOS14, outperforming the previous SOTA by 1.2%. For ActivityNet-1.3, despite its large scale and high diversity leading to lower absolute scores, our method attains an average mAP of 42.8%, exceeding the previous SOTA by 0.8%.
THUMOS14 & ActivityNet-1.3 under end-to-end setting. As shown in Table II, our proposed MambaTAD achieves SOTA performance across both datasets with lower memory usage, demonstrating its superior capability in temporal action detection. On THUMOS14, MambaTAD with VideoMAE-Huge as the backbone achieves an average mAP of 74.7%, outperforming the previous best method (AdaTAD with VideoMAE-Huge) by 0.4% while using less memory (2.5G less). Notably, even with the lighter VideoMAE-Large backbone, MambaTAD attains 74.3%, surpassing most existing approaches while maintaining a significantly lower memory footprint (9.3G vs. 19.2G for AdaTAD with VideoMAE-Huge). On ActivityNet-1.3, MambaTAD reaches an average mAP of 43.1% and 43.8% with the VideoMAE-Huge and VideoMAEv2-Giant [wang2023videomae] backbone, respectively. It not only significantly surpasses AdaTAD under the same backbone (+3.7%), but also outperforms the non-end-to-end performance of our MambaTAD with more expressive features InternVideo6B (+1.0%), demonstrating its efficiency and adaptability across different models and scales.
@@ IV-C Qualitative Evaluation
Statistical Visualization. Further, we present more specific visualizations of our results in Fig. 5, including the predicted action scores, and the regression outputs weighted by the action scores (as a weighted histogram). Our model effectively captures video sequences of varying lengths and adapts to action instances of different durations. Even in complex scenarios with multiple action instances, our approach consistently identifies action centers[zhang2022actionformer] and provides accurate boundary predictions.
@@ IV-D Robustness Across Varying Action Coverage and Length
Fig. 6 compares TAD methods using Coverage and Length metrics [alwassel2018diagnosing]. MambaTAD consistently performs well across coverage levels, excelling in the $0–4\%$ range for short actions and maintaining stability in the >>8% category, where others significantly decrease. This suggests strong long-range dependency modeling. For long actions (>>18ss), all methods struggle, but MambaTAD shows a more gradual performance drop, indicating better temporal robustness. These results confirm its effectiveness in handling varying action durations, driven by our DMBSS and global feature fusion head, which improve the modeling of complex, extended instances. Moreover, we conduct a detailed sensitivity and error analysis in Supp. B.
@@ IV-F Ablation Study
Impact of designs of DMBSS. This paper tackles two key challenges in applying Mamba to TAD: (1) temporal context decay in unidirectional processing and (2) diagonal conflicts from integrating two triangle matrices. We address these with bidirectional parameter sharing, a dual-branch approach, and diagonal masking. As shown in Table VIII, bidirectional parameter sharing alone reduces our method to DBM [chen2024video], while a dual-branch structure without sharing improves mAP to 73.3%, confirming its effectiveness in mitigating temporal decay. However, combining both leads to a slight accuracy drop due to diagonal conflicts. Resolving these conflicts restores accuracy, adding 0.9% (first line to fourth). Results further show that diagonal conflicts are minimal without parameter sharing, but masking significantly improves modeling performance. Overall, our approach ensures robust boundary representation by balancing bidirectional processing and conflict resolution.
Impact of the adapter. We conduct ablation studies on adapters using the VideoMAE-Large backbone, as shown in Table IX. First, employing the standard adapter[houlsby2019parameter] with the ActionFormer [zhang2022actionformer] detector as backbone results in an average mAP of 72.4%. Replacing the standard adapter with our state-space temporal adapter (SSTA) improves the average mAP by 0.9%. Similarly, substituting the ActionFormer detector with our detector also leads to a 0.9% performance gain. It is worth noting that, compared to Temporal-Informative Adapter (TIA), using our adapter alone or combining TIA with our detector results in a slight performance drop. We attribute this to two possible reasons: first, the reduction in parameter count and computational cost may have affected the model’s representational capacity; second, there may be an implicit architectural gap when training Mamba-based components alongside non-Mamba structures. In contrast, when both the adapter and detector are Mamba-based, the performance improves significantly. In the unified end-to-end framework, our end-to-end Mamba achieves a 1.9% improvement over the baseline, demonstrating the strong compatibility of DMBSS with the detector in an end-to-end setting and highlighting the effectiveness of SSTA.
@@
Mamba2 [dao2024transformers] Mamba2 83.8 72.6 42.8 67.9
Hydra [hwang2024hydra] Hydra 84.4 72.2 43.3 68.1
VideoMambaPro [lu2024videomambapro] VideoMambaPro 86.6 76.6 50.4 72.3
TimeMamba [chen2024video] DBM 86.5 75.9 49.9 72.7
Vision Mamba [DBLP:conf/icml/ZhuL0W0W24] ViM 86.9 76.9 52.3 73.0
CausalTAD [liu2024harnessing] DBM 87.4 77.6 50.8 73.1
@@ IV-G Feature Analysis
Figure 8: Intermediate feature visualization via t-SNE [van2008visualizing]. (a) the baseline method using the standard Mamba; (b) our proposed MambaTAD
@@ A Implementation details
We use single-stream pre-extracted InternVideo-6B features [wang2024internvideo2] and conduct fair comparisons with I3D [carreira2017quo] (THUMOS14) and R(2+1)D [alwassel2021tsp] (ActivityNet-1.3) features. For end-to-end training, we adopt VideoMAE [tong2022videomae] as the backbone, keeping all but the adapter frozen, with learning rates of 1e41e-4 (THUMOS14), 1e51e-5 (ActivityNet-1.3), and 4e44e-4 (MultiThumos). Our detection framework includes two convolutional embedding layers (except MultiThumos) and is optimized with AdamW [loshchilov2017decoupled]. The DMBSS kernel size is 4. Projection layer output dimensions are 512 (THUMOS14, MultiThumos, FineAction) and 256 (ActivityNet-1.3, HACS). We use 7 DMBSS projection layers for THUMOS14, ActivityNet-1.3, and HACS; 9 for FineAction; and 5 for MultiThumos, which also includes 3 convolutional embedding layers for multi-label scenarios. The global feature fusion head partitions features for classification and regression, following [shi2023tridet] for hyperparameters and bins. Our results are obtained by averaging over three independent runs. All experiments are conducted using PyTorch 2.0 and run on a single NVIDIA RTX A5000 GPU, except for the end-to-end setting on the ActivityNet-1.3 dataset, which utilizes 4 A5000 or A40 GPUs.

### v2

@@ III-D Building Projection Layers
The blocks designed in previous TAD networks [91, 62, 80] mainly follow the $Norm\rightarrow Attention/CNN\rightarrow Norm\rightarrow MLP$ flow in each pyramid layer. To address challenges related to computational complexity [97], feature discrimination [62, 80], and receptive field limitations [53, 17], we employ Mamba [15] to model global dependencies. However, in Section III-C, we analyze that the vanilla Mamba’s causality mechanism leads to a loss of temporal information, making it unsuitable for the TAD task. This underscores the potential of the proposed DMBSS block in effectively capturing action features for improved localization. The experimental results in Table VII in Section IV-F further validate our analysis.
The DMBSS projection layer. In this section, we construct projection layers utilizing the DMBSS discussed previously. As shown in Fig. 2, each projection layer consists of a DMBSS and a Max pooling layer [69], because we aim to preserve the benefits of long sequence modeling while enhancing the discriminability of temporal features, all within the linear computational complexity. The two blocks are given by:
@@ III-E Global Feature Fusion Head
The detection head plays a pivotal role in TAD, as it is directly linked to the final prediction outcome. Inspired by [85, 99], it is evident that the visual state-space model excels in tasks characterized by long sequences. Consequently, we innovatively concatenate features from different levels of the pyramid sequentially to create an extended sequence:
@@
where $\mathbf{F}\in\mathbb{R}^{b\times c\times(s+{s\over 2}+\ldots+{s\over 2^{l-1}})}$, and $l$ is the layer of pyramid. To address the lack of global awareness caused by hierarchical feature structure [39, 43], a global feature fusion mechanism is designed to ensure the full use of long-range information across different layers and improve regression and classification accuracy. To this end, we propose the residual global level detection head to adapt DMBSS for action detection. As shown in Fig. 3, we first use the LayerNorm (LN) followed by the DMBSS to capture the temporal long-range dependencies. Then, a residual connection is utilized for easier optimization. Lastly, we use a successive LN to ensure stable feature distribution. The final output can be mathematically represented as follows:
@@
By concatenating features from different pyramid levels into an extended sequence, our design enables the detection head to access temporal information across multiple granularities simultaneously. Traditional hierarchical feature pyramid approaches [39, 91] process each level independently, which limits the receptive field and reduces global context. In contrast, the extended sequence allows DMBSS to operate over a longer, contiguous feature stream, capturing temporal long-range dependencies across all pyramid layers. The residual global feature fusion further stabilizes the feature distribution and facilitates optimization. Together, this design enhances global awareness and improves both classification and boundary regression accuracy in temporal action detection.
@@ III-F State-Space Temporal Adapter
In Sections III-D and III-E, we construct the action detector with projection layers and a global feature fusion head. End-to-end TAD demands strong temporal modeling to bridge low-level encoding and high-level action detection. However, fine-tuning the entire backbone is costly [45] and may hinder generalization. On the other hand, it remains unexplored whether Mamba can serve as an adapter to bridge large pre-trained foundational models[71, 73, 74] and downstream video understanding tasks. To address these, we propose SSTA, a lightweight yet expressive mechanism for refining temporal representations efficiently.
The standard adapter [22] is widely used in PEFT but focuses on the channel dimension, limiting its ability to capture temporal dependencies. AdaTAD [45] addresses this with the Temporal-Informative Adapter (TIA), combining temporal depth-wise convolution and a fully connected layer. However, depth-wise convolution constrains the receptive field [53, 11], while the fully connected layer increases trainable parameters. To overcome these challenges, we propose the State-Space Temporal Adapter (SSTA), as illustrated in Fig. 2. SSTA integrates the DMBSS to enhance temporal modeling. This design enables bidirectional feature interaction while maintaining controlled diagonal masking, effectively capturing both long-range dependencies and local discriminative patterns. The formulation of SSTA is given by:
@@
where $\mathbf{x},\mathbf{x}^{\prime\prime}\in\mathbb{R}^{t\times h\times w\times d}$ denote input and output features, $t$ represents the number of frames, $h$ and $w$ correspond to the spatial dimensions of the video, and $d$ is the hidden feature dimensionality. The function $DMBSS(\cdot)$ represents our proposed block. Following AdaTAD[45], we employ GELU [21] as the non-linear activation function, denoted as $\theta(\cdot)$. The transformation matrices $\mathbf{W}^{T}_{down}\in\mathbb{R}^{d\times\frac{d}{\lambda}}$ and $\mathbf{W}^{T}_{up}\in\mathbb{R}^{\frac{d}{\lambda}\times d}$ are trainable projection weights, where $\lambda$ is a scalable factor which controls intermediate dimensions. $\lambda>1$ is enforced to ensure parameter efficiency and reduction. One residual connection is used for better convergence.
@@ IV-A Experimental Setup
Dataset. We experiment on five challenging datasets: THUMOS14 [26], ActivityNet-1.3 [3], MultiThumos[84], HACS-Segment [95], and FineAction [48]. The THUMOS14 dataset includes 20 sports action categories, with 200 training and 213 testing untrimmed videos, covering 3,007 and 3,358 action instances. ActivityNet-1.3 features 200 categories across approximately 20K videos and 600 hours, divided into 10K training, 4,926 validation, and 5,044 test videos. MultiThumos introduces more complex scenarios, presenting a greater challenge for TAD. It extends the THUMOS14 dataset with dense, multilabel, frame-level action annotations spanning 30 hours across 400 videos. The dataset comprises 38,690 annotations covering 65 action classes, with an average of 1.5 labels per frame and 10.5 action classes per video, significantly increasing the task’s difficulty. HACS-Segment has 200 categories, with 37.6K training and 5,981 test videos. FineAction offers 103K temporal instances across 106 fine-grained actions in 17K untrimmed videos, with 8,440 training, 4,174 validation, and 4,118 testing. Only training and validation annotations are available. We adhere to standard protocols for training and evaluation.
@@ IV-B Comparison With the State of the Art Methods
THUMOS14 & ActivityNet-1.3 with off-the-shelf features. Firstly, we compare MambaTAD with leading TAD methods on THUMOS14 and ActivityNet-1.3, as shown in Table I. Typically, TAD models achieve better performance with advanced backbones and optical flow, as evidenced by the second part of the table. Our evaluation encompasses a wide range of models, including CNNs, GCNs, Transformer-based, Diffusion-based, and Mamba-based architectures. For a fair comparison, we provide the results of MambaTAD using traditional features in the first part, highlighting its strong adaptability and compatibility. Specifically, MambaTAD surpasses the previous SOTA by 0.5% on THUMOS14 with I3D [4] features and 1.2% on ActivityNet-1.3 using R(2+1)D [1] features. In the second part, MambaTAD achieves an average mAP of 73.9% on THUMOS14, outperforming the previous SOTA by 1.2%. For ActivityNet-1.3, despite its large scale and high diversity leading to lower absolute scores, our method attains an average mAP of 42.8%, exceeding the previous SOTA by 0.8%.
THUMOS14 & ActivityNet-1.3 under end-to-end setting. As shown in Table II, our proposed MambaTAD achieves SOTA performance across both datasets with lower memory usage, demonstrating its superior capability in temporal action detection. On THUMOS14, MambaTAD with VideoMAE-Huge as the backbone achieves an average mAP of 74.7%, outperforming the previous best method (AdaTAD with VideoMAE-Huge) by 0.4% while using less memory (2.5G less). Notably, even with the lighter VideoMAE-Large backbone, MambaTAD attains 74.3%, surpassing most existing approaches while maintaining a significantly lower memory footprint (9.3G vs. 19.2G for AdaTAD with VideoMAE-Huge). On ActivityNet-1.3, MambaTAD reaches an average mAP of 43.1% and 43.8% with the VideoMAE-Huge and VideoMAEv2-Giant [73] backbone, respectively. It not only significantly surpasses AdaTAD under the same backbone (+3.7%), but also outperforms the non-end-to-end performance of our MambaTAD with more expressive features InternVideo6B (+1.0%), demonstrating its efficiency and adaptability across different models and scales.
@@ IV-C Qualitative Evaluation
Statistical Visualization. Further, we present more specific visualizations of our results in Fig. 5, including the predicted action scores, and the regression outputs weighted by the action scores (as a weighted histogram). Our model effectively captures video sequences of varying lengths and adapts to action instances of different durations. Even in complex scenarios with multiple action instances, our approach consistently identifies action centers[91] and provides accurate boundary predictions.
@@ IV-D Robustness Across Varying Action Coverage and Length
Fig. 6 compares TAD methods using Coverage and Length metrics [2]. MambaTAD consistently performs well across coverage levels, excelling in the $0–4\%$ range for short actions and maintaining stability in the >>8% category, where others significantly decrease. This suggests strong long-range dependency modeling. For long actions (>>18ss), all methods struggle, but MambaTAD shows a more gradual performance drop, indicating better temporal robustness. These results confirm its effectiveness in handling varying action durations, driven by our DMBSS and global feature fusion head, which improve the modeling of complex, extended instances. Moreover, we conduct a detailed sensitivity and error analysis in Supp. B.
@@ IV-F Ablation Study
Impact of designs of DMBSS. This paper tackles two key challenges in applying Mamba to TAD: (1) temporal context decay in unidirectional processing and (2) diagonal conflicts from integrating two triangle matrices. We address these with bidirectional parameter sharing, a dual-branch approach, and diagonal masking. As shown in Table VIII, bidirectional parameter sharing alone reduces our method to DBM [6], while a dual-branch structure without sharing improves mAP to 73.3%, confirming its effectiveness in mitigating temporal decay. However, combining both leads to a slight accuracy drop due to diagonal conflicts. Resolving these conflicts restores accuracy, adding 0.9% (first line to fourth). Results further show that diagonal conflicts are minimal without parameter sharing, but masking significantly improves modeling performance. Overall, our approach ensures robust boundary representation by balancing bidirectional processing and conflict resolution.
Impact of the adapter. We conduct ablation studies on adapters using the VideoMAE-Large backbone, as shown in Table IX. First, employing the standard adapter[22] with the ActionFormer [91] detector as backbone results in an average mAP of 72.4%. Replacing the standard adapter with our state-space temporal adapter (SSTA) improves the average mAP by 0.9%. Similarly, substituting the ActionFormer detector with our detector also leads to a 0.9% performance gain. It is worth noting that, compared to Temporal-Informative Adapter (TIA), using our adapter alone or combining TIA with our detector results in a slight performance drop. We attribute this to two possible reasons: first, the reduction in parameter count and computational cost may have affected the model’s representational capacity; second, there may be an implicit architectural gap when training Mamba-based components alongside non-Mamba structures. In contrast, when both the adapter and detector are Mamba-based, the performance improves significantly. In the unified end-to-end framework, our end-to-end Mamba achieves a 1.9% improvement over the baseline, demonstrating the strong compatibility of DMBSS with the detector in an end-to-end setting and highlighting the effectiveness of SSTA.
@@
Mamba2 [10] Mamba2 83.8 72.6 42.8 67.9
Hydra [24] Hydra 84.4 72.2 43.3 68.1
VideoMambaPro [51] VideoMambaPro 86.6 76.6 50.4 72.3
TimeMamba [6] DBM 86.5 75.9 49.9 72.7
Vision Mamba [97] ViM 86.9 76.9 52.3 73.0
CausalTAD [44] DBM 87.4 77.6 50.8 73.1
@@ IV-G Feature Analysis
Figure 8: Intermediate feature visualization via t-SNE [72]. (a) the baseline method using the standard Mamba; (b) our proposed MambaTAD
@@ A Implementation details
We use single-stream pre-extracted InternVideo-6B features [74] and conduct fair comparisons with I3D [4] (THUMOS14) and R(2+1)D [1] (ActivityNet-1.3) features. For end-to-end training, we adopt VideoMAE [71] as the backbone, keeping all but the adapter frozen, with learning rates of 1e41e-4 (THUMOS14), 1e51e-5 (ActivityNet-1.3), and 4e44e-4 (MultiThumos). Our detection framework includes two convolutional embedding layers (except MultiThumos) and is optimized with AdamW [50]. The DMBSS kernel size is 4. Projection layer output dimensions are 512 (THUMOS14, MultiThumos, FineAction) and 256 (ActivityNet-1.3, HACS). We use 7 DMBSS projection layers for THUMOS14, ActivityNet-1.3, and HACS; 9 for FineAction; and 5 for MultiThumos, which also includes 3 convolutional embedding layers for multi-label scenarios. The global feature fusion head partitions features for classification and regression, following [62] for hyperparameters and bins. Our results are obtained by averaging over three independent runs. All experiments are conducted using PyTorch 2.0 and run on a single NVIDIA RTX A5000 GPU, except for the end-to-end setting on the ActivityNet-1.3 dataset, which utilizes 4 A5000 or A40 GPUs.